\noindent



\begin{thebibliography}{00}

%% \bibitem[Author(year)]{label}
%% Text of bibliographic item

%\bibitem[ ()]{}
\bibitem{akbary2011constructing} A. Akbary, D. Ghioca and Q. Wang, On
 constructing permutations of finite fields,  Finite Fields  Appl. 17 (2011) 51--67.
%\bibitem{berlekamp1967on} E.R. Berlekamp, H. Rumsey, G.Solomon, On the solution of algebraic equations
%over finite fields, Inf. Control 10 (6) (1967) 553--564.
%%
%\bibitem{cepak2017permutation} N.Cepak, P. Charpin, E. Pasalic, Permutation via linear translators,
%Finite Fields Appl. 45 (2017) 19--42.
%%\bibitem{CK09} P. Charpin, G. Kyureghyan, When does $F(X)+ \Tr(H(X))$ permute $\gf_{p^n}$?, Finite Fields Appl. 15 (5), 615--632 (2009).
%%
%%\bibitem{CK08} P. Charpin, G. Kyureghyan, On a class of permutation polynomials over $\gf_{2^n}$, in: SETA 2008, in: Lecture Notes in Comput.
%%Sci., vol. 5203, Springer-Verlag, 368--376, 2008.
%%
\bibitem{coulter2002compositional} R. S. Coulter and M. Henderson, The compositional inverse of a class of permutation polynomials over a finite field, Bulletin of the Australian Mathematical Society, 65(3) (2002) 521--526.

\bibitem{ding2013cyclic} C. Ding, Cyclic codes from some monomials and trinomials, SIAM J. Discrete Math., 27 (2013) 1977--1994.
%%
%\bibitem{ding2013optimal} C. Ding, T. Helleseth, Optimal ternary cyclic codes from monomials, IEEE
%Trans. Inf. Theory 59 (2013) 5898--5904.
%%
\bibitem{ding2014binary} C. Ding, Z. Zhou, Binary cyclic codes from explicit polynomials over $GF(2^m)$, Discrete Math., 321 (2014) 76--89.
%%
%%
%%
%%
%%
%%
%%\bibitem{DQWYY15} C. Ding, L. Qu, Q. Wang, J. Yuan, P. Yuan, Permutation trinomials over finite
%%fields with even characteristic, SIAM J. Discrete Math. 29, 79--92 (2015).
%%
%%
%%\bibitem{DXY}  C. Ding, Q. Xiang, J. Yuan, P. Yuan, Explicit classes of permutation polynomials over GF($3^{3m}$),
%%Sciences in China Ser. A 53, 639--647 (2009).
%%
\bibitem{ding2006family} C. Ding and J. Yuan, A family of skew Hadamard difference sets,
J. Comb. Theory, Ser. A 113 (2006) 1526--1535.
%%
%%\bibitem{GS18}R. Gupta, R.K.Sharma, Further results on permutation polynomials of the form $(x^{p^m}-x+\delta)^s+x $ over $\gf_{p^{2m}},$ Finite Fields Appl. 50, 196--208(2018).
%%
%\bibitem{helleset2003new} T. Helleset, V. Zinoviev,: New Kloostermans sums identities over $\gf_{2^m}$ for all $m$, Finite Fields Appl. 9(2) (2003) 187--193.
%%
%%
%%
%\bibitem{Hou2015survey} X. Hou, A survey of permutation binomials and trinomials over finite fields, in:
%Proceedings of the 11th International Conference on Finite Fields and Their Applications, Contemp. Math., Magdeburg, Germany, July 2013, 632 AMS (2015)
%177--191.
%%\bibitem{Hou152} X. Hou, Permutation polynomials over finite fields - A survey of recent advances,
%%Finite Fields Appl. 32, 82--119 (2015).
%%
%%\bibitem{Ky10} G. Kyureghyan, Constructing permutations of finite fields via linear translators,
%%J. Combin. Theory Ser. A 118, 1052--1061 (2010).
\bibitem{laigle2007permutation} Y. Laigle-Chapuy, Permutation polynomial and applications to coding theory,
Finite Fields Appl. 13 (2007) 58--70.
%%\bibitem{LHT13} N. Li, T. Helleseth, X. Tang, Further results on a class of permutation polynomials over finite fields, Finite Fields Appl. 22, 16--23 (2013).
%%
%%
%%\bibitem{LP97} J.B. Lee, Y.H. Park, Some permutation trinomials over finite fields, Acta Math.
%%Sci., 17, 250--254 (1997).
%%
%%\bibitem{LQC15} K. Li, L. Qu, X. Chen, New classes of permutation binomials and permutation
%%trinomials over Finite Fields, Finite Fields Appl. 43, 69--85 (2017).
%%
%%
\bibitem{li2023several} G. Li, X. Cao, Several classes of permutation polynomials over the finite field $\gf_{p^{2m}}.$ Finite Fields Appl. 89 (2023) 102197.


\bibitem{li2024several} G. Li, X. Cao, Several classes of permutation polynomials  based on the AGW criterion over the finite field $\gf_{p^{2m}}.$ Finite Fields Appl. 95 (2024) 102378.


\bibitem{li2019compositional} K. Li, L. Qu, Q. Wang, Compositional inverses of permutation polynomials of the form $x^rh(x^s)$ over finite fields, Cryptogr. Commun., 11 (2019) 279--298.


\bibitem{li2018permutation} L. Li, S. Wang, C. Li, N. Li, X. Zeng, Permutation polynomials
$(x^{p^m}-x+\delta)^{s_1}+(x^{p^m}-x+\delta)^{s_2}+x$ over $\gf_{p^n},$ Finite Fields Appl. 51 (2018) 31--61.
%%
\bibitem{lidl1997finite} R. Lidl, H. Niederreiter,
Finite Fields, second ed., Encyclopedia of Mathematics and its
Applications, vol. 20, Cambridge University Press, Cambridge, 1997.

\bibitem{lidl1986introduction} R. Lidl, H. Niederreiter,
Introduction to Finite Fields and Their Applications,  Cambridge University Press, Cambridge, 1986.
%%
%%\bibitem{Ma11} J. E. Marcos, Specific permutation polynomials over finite fields,
%%Finite Fields Appl. 17, 105--112(2011).
%%
%%\bibitem{MZFG15} J. Ma, T. Zhang, T. Feng, and G. Ge, Some new results on permutation polynomials over finite fields, Des. Codes Cryptogr. (2016). DOI: 10.1007/s10623-016-0236-1.
%%
%%
\bibitem{mull1993permutation} G. L. Mullen, Permutation polynomials over finite fields, In: Proc.
Conf. Finite Fields and Their Applications, Lecture Notes in Pure and Applied
Mathematics, vol. 141, Marcel Dekker, (1993) 131--151.
%%
%%\bibitem{MullWang} G. L. Mullen, Q. Wang, Permutation polynomials in one variable, pp.215--229, in: G. L. Mullen, D. Panario (Eds.), Handbook of Finite Fields, CRC Press,2013.
%%

\bibitem{niu2021finding} T. Niu, K. Li, L. Qu, Q. Wang, Finding compositional inverses of permutations form the AGW criterion, IEEE Trans. Inf. Theory, 67 (2021) 4975--4985.


%%\bibitem{PL01} Y. H. Park, J. B. Lee, Permutation polynomials and group
%%permutation polynomials, Bull. Austral. Math. Soc. 63,
%%67--74 (2001).
%%
%%
\bibitem{reis2021permutation} L. Reis, Q. Wang, Permutation polynomials from a linearized decomposition,  arXiv:2104.13234v1 (2021).

\bibitem{reis2024constrcuting} L. Reis, Q. Wang, Constructing permutation polynomials from permutation polynomials of subfields, Finite Fields Appl., 96 (2024) 102415.



\bibitem{rivest1978method} R. L. Rivest, A. Shamir, and L. M. Adelman, A method for obtaining digital signatures and public-key cryptosystems,
Comm. ACM 21 (1978) 120--126.
%%
\bibitem{schwenk1998public}
J. Schwenk and K. Huber,
Public key encryption and digital signatures based on permutation polynomials,
Electronic Letters, 34 (1998) 759--760.

%%
\bibitem{tu2015two}Z.Tu, X. Zeng, X. Jiang, Two classes of permutation polynomials having the
form $(x^{2^m}+x+\delta)^s+x, $  Finite Fields Appl. 31 (2015) 12--24.

\bibitem{tuxanidy2014inverses} A. Tuxanidy, Q. Wang, On the inverses of some classes of permutations of finite fields, Finite Fields and Their Applications, 28 (2014) 244--281.

\bibitem{tuxanidy2017compositional} A. Tuxanidy, Q. Wang, Compositional inverses and complete mappings over finite fields, Discrete Applied Mathematics, 217 (2017) 318--329.

%\bibitem{wang2007cyclotomic} Q. Wang, Cyclotomic mapping permutation polynomials over finite fields, Sequences, Subsequences, Consequences, 4893 (2007) 119--128.

\bibitem{wang2017note} Q. Wang, A note on inverses of cyclotomic mapping permutation polynomials over finite fields, Finite Fields Appl., 45 (2017) 422--427.



\bibitem{baofengwu2013compositional} B. Wu, The compositional inverses of linearized permutation binomials over finite fields, arXiv:1311.2154v1 (2013).
 
 
 \bibitem{wu2014compositional} B. Wu, The compositional inverse of a class of linearized permutation polynomials over $\gf_{2^n}$, $n$ odd,  Finite Fields Appl.,  29 (2014) 34--48.
 
%\bibitem{wu2013compositional} B. Wu, Z. Liu, The compositional inverse of a class of bilinear permutation polynomials over finite fields of characteristic $2$, Finite Fields Appl. 24 (2013) 136--147.
 
 
%\bibitem{Baofengwu2013linearized} B. Wu, Z. Liu, Linearized polynomials over finite fields revisited, Finite Fields Appl., 22 (2013) 79--100. 
%%\bibitem{TZLH15}Z.Tu, X. Zeng, C. Li, T. Helleseth, Permutation polynomials of the from $(x^{p^m}-x+\delta)^s+L(x) $ over the finite field $\gf_{p^{2m}}$ of odd characteristic,  Finite Fields Appl. 34, 20--35 (2015).
%%
%%\bibitem{Wa07} Q. Wang, Cyclotomic mapping permutation polynomials over finite
%%fields, Sequences, subsequences, and Consequences (International
%%Workshop, SSC 2007, Los Angeles, CA, USA, May 31 - June 2, 2007),
%%Lecture Notes in Comput. Sci. 4893, 119--128.
%%
%%\bibitem{WWL17}L. Wang, B. Wu, Z. Liu, Further results on permutation polynomials
%%of the form $(x^{p}-x+\delta)^s+x $ over  $\gf_{p^{2m}}$, Finite Fields Appl. 44, 92--112  (2017).
%%
%%\bibitem{WW18}L. Wang, B. Wu,  General constructions of  permutation polynomials
%%of the form $(x^{2^m}+x+\delta)^{i(2^m-1)+1}+x $ over  $\gf_{2^{2m}}$, Finite Fields Appl. 52, 137--155 (2018).
%%
\bibitem{wu2022some} D. Wu, P. Yuan, Some classes of permutation polynomials of the form
$b(x^q+ax+\delta)^{1+i(q^2-1)/d}+c(x^q+ax+\delta)^{1+j(q^2-1)/d}+L(x)$ over $\gf_{q^{2}}, $ Applicable Algebra in Engineering, Communication and Computing. 33 (2022) 135-149.

\bibitem{wu2023some} D. Wu, P. Yuan, Some new results on permutation polynomials of the form $b(x^q+ax+\delta)^s-ax$ over $\gf_{q^2},$ Finite Fields Appl. 93 (2024) 102329.
%%
\bibitem{xu2022several}G. Xu, G. Luo, X. Cao, Several classes of permutation polynomials of the
form $(x^{p^m}-x+\delta)^{s}+x $ over  $\gf_{p^{2m}}$, Finite Fields Appl. 79 (2022) 102001. 


%%
%\bibitem{yuan2007four} J. Yuan, C. Ding, Four classes of permutation polynomials over $\gf_{2^m}$, Finite Fields Appl. 13 (4) (2007) 869--876.
%%%
%%%
%\bibitem{yuan2008permutation} J. Yuan, C. Ding, H. Wang, J. Pieprzyk, Permutation polynomials of the form $(x^p-x+\delta)^s+L(x),$ Finite Fields Appl. 14(2) (2008) 482--493.
%%
%%
%%
%%
%%
%%\bibitem{YD07} J. Yuan, C. Ding, Four classes of permutation polynomials of $\gf_{2^m}, $ Finite Fields Appl. 13(4), 869--876 (2007).
%%
%%\bibitem{YDWP08} J. Yuan, C. Ding, H. Wang, J. Pieprzyk,  Permutation polynomials of
%%  the form $(x^p-x+\delta)^s+L(x),$ Finite Fields Appl. 14(2),  482--493 (2008).
%%
\bibitem{yuan2022compositional}
P. Yuan, Compositional inverses of AGW-PPs, Advances in Mathematics of Communications, 16(4)
 (2022) 1185--1195.
 
 \bibitem{yuan2022agwpps} P. Yuan, Compositional inverses of AGW-PPs, Adv. Math. Comm. 2022, 16(4): 1185--1195.
 \bibitem{yuan2022local} P. Yuan, Local method for compositional inverses of permutation polynomials, Commun. Algebra, 52(7) (2024) 3070--3080.

%\bibitem{yuan2011permutation} P. Yuan, C. Ding, Permutation polynomials over finite fields from a powerful lemma, Finite Fields Appl. 17 (2011) 560--574.
%%
%%\bibitem{YD14} P. Yuan, C. Ding, Further results on permutation polynomials over finite fields, Finite Fields Appl. 35, 215--230 (2015).
%%
\bibitem{yuan2015permutation} P. Yuan, Y. Zheng  , Permutation polynomials from piecewise funtions, Finite Fields Appl. 35 (2015) 215--230.
%%

\bibitem{zheng2019inverses} Y. Zheng, Q. Wang and W. Wei, On inverses of permutation polynomials of small degree over finite fields, IEEE Trans. Inf. Theory, 66(2) (2019), 914--922.


\bibitem{zheng2019constructions} D. Zheng, M. Yuan, N. Li, L. Hu, X. Zeng,
 Constructions of involutions over finite fields,  IEEE Trans. Inf. Theory, 65(12) (2019), 7876--7883.
 
%%\bibitem{ZZH} X. Zeng, X. Zhu and L. Hu, Two new permutation polynomials
%%with the form $(x^{2^k}+x+\delta)^s+x$ over $\gf_{2^n}$,
%%Applicable Algebra in Engineering, Communication and Computing
%%21, 145--150 (2010).
\bibitem{zeng2017permutation} X. Zeng, X. Zhu, N. Li, X. Liu, Permutation polynomials over $\gf_{2^n}$
of the form $(x^{2^i}+x+\delta)^{s_1}+(x^{2^i}+x+\delta)^{s_2}+x$, Finite Fields Appl. 47, 256-268 (2017).
%%
%%\bibitem{ZH} Z. Zha, L. Hu, Two classes of permutation polynomials over finite fields, Finite Fields Appl. 18,  781--790 (2012).
%%\bibitem{ZH16} Z. Zha, L. Hu, Some classes of permuation polynomials of the form $(x^{p^m}-x+\delta)^s+L(x)$ over $\gf_{p^{2m}}$, Finite Fields Appl. 40,  43--52 (2016).
%%
%%
%%\bibitem{ZC17} D. Zheng, Z. Chen, More classes of permutation polynomials of the form
%%$(x^{2^m}+x+\delta)^s+L(x),$ Appl. Algebra Eng. Commun. Comput. 28(3),215-223 (2017).
%%
%%\bibitem{ZYY19} D. Zheng, M. Yuan, L. Yuan, Two types of permutation polynomials with
%%special forms, Finite Fields Appl. 56, 1--16 (2019).
%%
%%
%%\bibitem{Zi08} M. E. Zieve, Some families of permutation polynomials over finite
%%fields. Internat. J. Number Theory 4, 851--857 (2008).
%%
%%
%%\bibitem{Zi09}
%%M. E. Zieve, On some permutation polynomials over $\gf_q$ of the
%%form $x^rh(x^{(q-1)/d})$, Proc. Amer. Math. Soc. 137, 209--216 (2009).
%%
%%\bibitem{Zi10} M. E. Zieve, Classes of permutaton polynomials based
%%on cyclotomy and an additive analogue, pp. 355--361, in: Additive Number Theory,
%%Springer, 2010.
%%
%%\bibitem{Zi13} M.E. Zieve, Permutation polynomials on $\gf_q$ induced from Redei function bijections on subgroups of $\gf_q^\ast$, arXiv:1310.0776., 2013.

\end{thebibliography}

\documentclass[preprint,12pt,authoryear]{elsarticle}

\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{enumerate}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{enumitem}
\usepackage{booktabs}
\usepackage{caption}
\usepackage{natbib}
\setcitestyle{numbers,square}
\usepackage{multirow}
\usepackage{tabularx}

\newcommand{\Tr}{{\rm Tr}}
\newcommand{\Ker}{{\rm Ker}}
\newcommand{\gf}{ {{\mathbb F}} }
\newtheorem{lemma}{Lemma}[section]
\newtheorem{proposition}{Proposition}[section]
\newtheorem{theorem}{Theorem}[section]
\newtheorem{remark}[theorem]{Remark}

\journal{Finite Field and Their Applications}

\begin{document}

\begin{frontmatter}



\title{The compositional inverses of permutation polynomials of the form $\sum_{i=1}^kb_i(x^{p^m}+x+\delta)^{s_i}-x$ over $\gf_{p^{2m}}$}


\author[wuaddress]{Danyao Wu\corref{mycorrespondingauthor}}
\cortext[mycorrespondingauthor]{Corresponding author}
\ead{wudanyao111@163.com}

\author[yuanaddress]{Pingzhi Yuan}
\ead{yuanpz@scnu.edu.cn}
\author[guanaddress]{Huanhuan Guan} 
\ead{guan1110h@163.com}

\author[liaddress]{Juan Li}
\ead{41170208@qq.com}

\address[wuaddress]{School of Computer Science and Technology,
	Dongguan University of Technology, Dongguan 523808, China}
\address[yuanaddress]{School of Mathematics, South China Normal University, Guangzhou 510631, China}
\address[guanaddress]{School of Mathematics and Statistics, Guizhou University of Finance and Economics, Guiyang 550025, China}
\address[liaddress]{School of Mathematics, Jiaying University, Meizhou, 514015, China}

%
%
%
%
%
%

%

\begin{abstract}
In this paper, we present the compositional inverses of several classes permutation polynomials of the form $\sum_{i=1}^kb_i(x^{p^m}+x+\delta)^{s_i}-x$ over $\gf_{p^{2m}}$, where for $1\leq i \leq k,$ $s_i, m$ are positive integers, $b_i, \delta \in \gf_{p^{2m}},$  and $p$ is prime.

\end{abstract}



\begin{keyword}


Finite field \sep Compositional inverse \sep Permutation polynomial
\MSC 11C08 \sep 12E10
\end{keyword}

\end{frontmatter}


\section{Introduction}
\label{}
Let  $\gf_q$ be the finite field with $q$ elements,  where $q$ is a prime power, and
let $\gf_q[x]$
be the ring of polynomials in a single indeterminate $x$ over $\gf_q$. A polynomial
$f \in\gf_q[x]$ is called a {\em permutation polynomial} of $\gf_q$ if its
associated polynomial mapping $f: c\mapsto f(c)$ from $\gf_q$ to itself is bijective. The unique polynomial denoted by $f^{-1}(x)$ over $\gf_q$
such that $f(f^{-1}(x))\equiv f^{-1}(f(x)) \equiv x \pmod{x^q-x}$ is called the compositional inverse of $f(x).$ Furthermore,  $f(x)$ is called  an involution when $f^{-1}(x)=f(x).$

The study of permutation polynomials and their compositional inverses over finite
fields in terms of their coefficients is a classical and difficult subject which
attracts people's interest partially due to their wide applications in coding theory
\cite{ding2013cyclic,ding2014binary,laigle2007permutation},
cryptography \cite{rivest1978method,schwenk1998public}, combinatorial design theory \cite{ding2006family}, and other areas of mathematics and engineering \cite{lidl1997finite,lidl1986introduction,mull1993permutation}. For instance, in block ciphers,
a permutation polynomial is usually used as an S-box to build the confusion layer and the compositional inverse of S-box comes into picture while decrypting the cipher. Both the permutation polynomial and its compositional inverse are implemented. Therefore, the explicit and efficient permutation polynomial and its compositional inverse are desired for designers. Indeed, gaining 
a better understanding of permutation polynomials and their compositional inverses in explicit format is not only meaningful, but also important for these applications.%
In general, it is difficult to discover new classes of permutation polynomials and computing the coefficients of
the compositional inverse of a permutation polynomial seems to be even more difficult, except for several classical classes such as
monomials, linearized polynomials, Dickson polynomials.
Compositional inverses of several classes of permutation polynomials
in explicit or implicit forms have been investigated in recent years,
which have nice structure.
We refer the readers to \cite{coulter2002compositional,li2019compositional,niu2021finding,tuxanidy2014inverses,tuxanidy2017compositional,wang2017note,baofengwu2013compositional,wu2014compositional,yuan2022compositional,yuan2022local,zheng2019inverses,zheng2019constructions}
for more details.

In 2014, Tuxanidy and Wang \cite{tuxanidy2014inverses}  studied the compositional inverse of a class of permutation polynomials of the form $f(x)=h(\psi(x))\varphi(x)+g(\psi(x))$ over $\gf_{q^n},$ where $\varphi, \psi, \bar{\psi}$ are additive polynomials satisfying $\varphi \circ \psi= \bar{\psi} \circ \varphi$ and $\sharp\psi(\gf_{q^n})=\sharp\bar{\psi}(\gf_{q^n}).$ Here,  $g, h \in \gf_{q^n}[x]$ are polynomials with $h(\psi(\gf_{q^n}))\subseteq \gf_{q}^*.$  The inverse of $f(x)$ is related to the inverses $\bar{f}(x)$ and $\varphi|_{\ker(\psi)}$, where $\bar{f}(x)=h(x)\varphi(x)+\bar{\psi(g(x))}$ is a bijection from $\psi(\gf_{q^n})$ to $\bar{\psi}(\gf_{q^n}).$ 

In 2021, Reis and Wang \cite{reis2021permutation} investigated the permutation polynomial of the form $P_1(x)=g(\Tr_{q^n/q}(x))+k(\Tr_{q^n/q}(x))L_h(x)$ and its inverse over $\gf_{q^n}$,  where  $g(x) \in \gf_{q^n}[x], h(x)\in \gf_q[x],$ $k(x) \in \gf_q[x]$ with  $k(\gf_q)\subseteq \gf_q^*.$ Here,  $L_h(x)$ is the linearized $q$-associate of $h.$ The compositional inverse of $P_1(x)$ is related to the inverse of  another permutation polynomial $Q(x)$ over $\gf_{q}.$ Later, they \cite{reis2024constrcuting} refined their results by  studying the permutation polynomial of $P_2(x)$ and its inverse over $\gf_{q^n}.$ In this case, $P_2(x)$ is obtained from $P_1(x)$  by replacing $L_l(x)$ with $\Tr_{q^n/q}(x)$,  and $L_l(x)$ is the $q$-associated of $l(x)=(x^n-1)/(x^t-a)$ with $t|n$ and $ a^{n/t}=1.$ 

Many series of explicit  permutation polynomials of the form $\sum_{i=1}^kb_i(x^{p^m}+ax+\delta)^{s_i}-ax$ $(a^{q+1}=1)$ over $\gf_{q^2}$ \cite{li2023several,li2024several,li2018permutation,wu2022some,wu2023some,xu2022several,yuan2015permutation,zeng2017permutation}  are special cases of $f(x)$ in \cite[Theorem 1.2]{tuxanidy2014inverses} ( where $n=2,$ $h(x)=1$, $\varphi(x)=-ax,$ $\psi(x)=x^q+ax,$ $\bar{\psi}(x)=ax^q+x$, and $g(x)=\sum_{i=0}^k(x+\delta)^{s_i}$),  or  $P_1(x)$ in \cite[Theorem 3.2]{reis2021permutation} approximately (see \eqref{P(x)} and $n=2,$  $k(x)=1, $ $L_h(x)=-x$,  $g(x)=\sum_{i=0}^k(x+\delta)^{s_i}$). However, finding $
\bar{f}^{-1}(x)$ in \cite[Theorem 1.2]{tuxanidy2014inverses} and $Q^{-1}(x)$ in \cite[Theorem 3.2]{reis2021permutation} are still challenging. In this paper, we will explicitly determine  the compositional inverses of the permutation polynomials of the form $\sum_{i=1}^kb_i(x^{p^m}+x+\delta)^{s_i}-x$ over $\gf_{q^2}$ with different choices of the exponents $s_i.$

The remainder of this paper is organized as follows. Section 2  introduces  related results. Next, Section 3 presents the investigation of the compositional inverses of permutation polynomials in the form $\sum_{i=1}^kb_i(x^{p^m}+x+\delta)^{s_i}-x$ over $\gf_{p^{2m}}.$

\section{Auxiliary results }
In this section, we present some auxiliary results that will be needed in the
sequel.

\begin{lemma}\cite[Theorem 5.1 ]{akbary2011constructing}\label{th5.2} For any polynomial $g \in \gf_{q^n}[x], $ any  additive polynomials 
 $\varphi, \psi, \bar{\psi}$ satisfying $\varphi \circ \psi= \bar{\psi} \circ \varphi$ and $\sharp\psi(\gf_{q^n})=\sharp\bar{\psi}(\gf_{q^n}),$ and any polynomial $h \in \gf_{q^n}[x]$ such that  $h(\psi(\gf_{q^n}))\subseteq \gf_{q}^*.$ Then $$P(x)=h(\psi(x))\varphi(x)+g(\psi(x))$$ permutes $\gf_{q^n}$ if and only if 
 
 (i) $\ker(\varphi)\cap \ker(\psi)=\{0\},$ or, equivalently, $\varphi$ induces a bijection from $\ker(\psi)$ to $\ker(\bar{\psi});$   and 
 
 (ii) $\tau(x)=h(x)\varphi(x)+\bar{\psi}(g(x))$ is a bijection from $\psi(\gf_{q^n})$ to $\bar{\psi}(\gf_{q^n}).$
\end{lemma}
		
Even though $\varphi(x)\circ \psi(x)=\bar{\psi}(x) \circ \varphi(x)$ and $\sharp \psi(\gf_{q^n})=\sharp \bar{\psi}(\gf_{q^n})$ hold, we still may not  have a "nice enough" expression for   
$\Ker(\psi).$ 
To calculate the inverse of $f(x)$ in Lemma \ref{th5.2},   Tuxanidy and Wang \cite{tuxanidy2014inverses}  added the conditions $\sharp S_{\psi}=\sharp S_{\bar{\psi}}$ and $\ker(\varphi)\cap \psi(S_{\psi})=\{0\},$ 
where  $S_{\psi}=\{x-\psi(x)\mid x\in \gf_{q^n}\}$ and $S_{\bar{\psi}}=\{x-\bar{\psi}(x)\mid x\in \gf_{q^n}\}.$  Then they have the following result.

\begin{lemma}\cite[Theorem 1.2]{tuxanidy2014inverses}\label{th1.2}
	Using the same notation and assumptions of Lemma \ref{th5.2}, assume that $P(x)$ is a permutation of $\gf_{q^n}.$ Let $\tau^{-1}, $ $\varphi^{-1}|_{S_{\bar{\psi}}}$ induce the inverses of $\tau|_{\psi(\gf_{q^n})}$ and $\varphi|_{S_{\psi}},$ respectively. Then the compositional inverse of $P(x)$ on $\gf_{q^n}$ is given by 
	$$P^{-1}(x)=\tau^{-1}(\bar{\psi}(x))+\varphi^{-1}|_{S(\bar{\psi})}\left(\frac{x-\bar{\psi}(x)-g(\tau^{-1}(\bar{\psi}(x)))+\bar{\psi}(g(\tau^{-1}(\bar{\psi}(x))))}{h(\tau^{-1}(\bar{\psi}(x)))}\right).$$ 
	Furthermore, if $\varphi$ induces a bijection from $\psi(\gf_{q^n})$ to $\bar{\psi}(\gf_{q^n}),$ then $\varphi$ permutes $\gf_{q^n}$ and the compositional inverse of $P(x)$ on $\gf_{q^n}$ is given by 
		$$P^{-1}(x)=\varphi^{-1}\left(\frac{x-g(\tau^{-1}(\bar{\psi}(x)))}{h(\tau^{-1}(\bar{\psi}(x)))}\right).$$ 
 \end{lemma}

Now, taking $h(x)=1,$ $\varphi(x)=-x,$ $\psi(x)=\bar{\psi}(x)=\Tr_{q^2/q}(x),$ and $g(x)=\sum_{i=1}^kb_i(x+\delta)^{s_i}$ in Lemmas \ref{th5.2} and \ref{th1.2},  then we have the following lemma. 
\begin{lemma}\label{th1}
	Let $q$ be a prime power.
	For $i=1, 2, \cdots, d,  $ assume that $b_i, \delta \in \gf_{q^2}$ and $s_i$ are positive integers. 
	Then the polynomial  $$P(x)=\sum_{i=1}^{k}b_i(x^q+x+\delta)^{s_i}-x$$ permutes $\gf_{q^2}$ if and only if 
	\begin{equation}\label{letau}
		\tau(x)=\Tr_{q^2/q}(x) \circ g(x)-x
		\end{equation}
	permutes $\gf_q.$ Moreover, if $p^{-1}(x)$ exists over $\gf_{q^2},$ then 
	$$P^{-1}(x)=g(x)\circ \tau^{-1}(x)\circ \Tr_{q^2/q}(x)-x,$$ where $g(x)=\sum_{i=1}^kb_i(x+\delta)^{s_i}.$
	\end{lemma}
	
Lemma \ref{th1}	is crucial in this paper, which will be frequently used in  next section.

We list two results about the compositional inverses of linearized permutation polynomials at last. 
 
For a positive integer $m, $ 
define a sequence 
$$S_{-1}=0, S_0=1, S_i=b^{2^{i-1}}S_{i-1}+a^{2^{i-1}}S_{i-2},$$ 
where $1\leq i\leq m$ and $a, b \in \gf_{2^m}.$ Y. Zheng, Q. Wang and W. Wei \cite{zheng2019inverses} studied  the inverse of linearized polynomial of the form $x^4+bx^2+ax$ over $\gf_{2^m}.$
\begin{lemma}\label{421} \cite[Corollary 4]{zheng2019inverses}
Let $L(x)=x^4+bx^2+ax$, where $a ,b \in \gf_{2^m}$ and $m>1.$ Then $L(x)$ is a permutation polynomial over $\gf_{2^m}$ if and only if $S_m+aS_{m-2}^2=1.$ Moreover, if $L(x)$ permutes $\gf_{2^m},$ the inverse of $L(x)$ over $\gf_{2^m}$ is given by  
$$L^{-1}(x)=\sum_{i=0}^{m-1}(S_{m-2-i}^{2^{i+1}}+a^{1-2^{i+1}}S_i)x^{2^i}.$$
	
	\end{lemma}

\begin{lemma}\label{binomial}\cite[Theorem 2.1]{baofengwu2013compositional}
Let $L_r(x)=x^{q^r}-ax$, where $a\in \gf_{q^m}^*$, and $1\leq r\leq m-1$. Then $L_r(x)$ is a permutation polynomial over $\gf_{q^m}$ if and only if the norm $N_{q^m/q^d}(a)\neq 1,$ where $d=gcd(m ,r).$ In this case, its inverse on $\gf_{q^m}$ is 
$$L_r^{-1}(x)=\frac{N_{q^m/q^d}(a)}{1-N_{q^m/q^d}(a)}\sum_{i=0}^{m/d-1}a^{-\frac{q^{(i+1)r}-1}{q^r-1}}x^{q^{ir}}.$$
\end{lemma}

\section{Main results}

 We consider any $a\in \gf_{q^2}$ with $a^{q+1}=1.$ Let $\varepsilon \in \gf_{q^2}$ be  a primitive element of $\gf_{q^2}$ and $t$ be the positive integer with
$a=\varepsilon^{(q-1)t}$ and $0\leq t\leq q.$ 
Put $x=\varepsilon^ty,$ we have that 
\begin{align}\label{P(x)}
	\sum_{i=1}^{k}b_i(x^q+ax+\delta)^{s_i}-ax=&\,
	\sum_{i=1}^{k}b_i(x^q+\varepsilon^{(q-1)t}x+\delta)^{s_i}-\varepsilon^{(q-1)t}x\nonumber\\
	=&\,\sum_{i=1}^{k}b_i\left(\varepsilon^{qt}(y^q+y+\varepsilon^{-qt}\delta)\right)^{s_i}-\varepsilon^{qt}y\nonumber\\
	=&\, \varepsilon^{qt}\left(	\sum_{i=1}^{k}\bar{b}_i(y^q+y+\bar{\delta})^{s_i}-y\right),
\end{align}
Where $\bar{b}_i=b_i\varepsilon^{qt(s_i-1)}$ and $\bar{\delta}=\varepsilon^{-qt}\delta.$ Hence, it suffices to consider the compositional inverses of  the permutation polynomials of  the form 

$$P(x)=\sum_{i=1}^{k}b_i(x^q+x+\delta)^{s_i}-x$$ 
over $\gf_{q^2}.$

We  give a proposition at first. 
\begin{proposition}\label{th11}
	Let $q$ be a prime power.  For $1\leq i \leq k,$  assume that $b_i, \delta \in \gf_{q^2}$ and $s_i$ are positive integers.  Then the polynomial $ P_1(x)=\sum_{i=1}^{k}b_i(x^q+x+\delta)^{s_i}-x$ is a permutation polynomial over $\gf_{q^2}$ if and only if $P_2(x)=\sum_{i=1}^{k}b_i^q(x^q+x+\delta)^{qs_i}-x$ permutes $\gf_{q^2}$. Moreover, if $P_1(x)$ is a permutation polynomial over $\gf_{q^2}$, then   $$P_1^{-1}(x)+x=(P_2^{-1}(x)+x)^q,$$
	where $P_1^{-1}(x)$ and $P_2^{-1}(x)$ are the compositional inverses of $P_1(x)$ and $P_2(x)$, respectively.
\end{proposition}
\begin{proof}
	Since $P_2(x^q)=\sum_{i=1}^{k}b_i^q(x^q+x+\delta)^{qs_i}-x^q=P_1(x)^q,$ we draw the first conclusion  immediately.  
	
	Moreover, if $P_1(x)$ permutes $\gf_{q^{2}}$, then $P_2(x)$ permutes $\gf_{q^2}.$ It follows from  Lemma \ref{th1}  that 
	$$P_1^{-1}(x)=\left(\sum_{i=1}^{k}b_i(x+\delta)^{s_i}\right)\circ \tau^{-1}(x)\circ  \Tr_{q^2/q}(x)-x$$ and $$P_2^{-1}(x)=\left(\sum_{i=1}^{k}b_i^q(x+\delta)^{qs_i}\right)\circ  \tau^{-1}(x)\circ  \Tr_{q^2/q}(x)-x,$$ where $\tau(x)=\Tr_{q^2/q}(x) \circ \left(\sum_{i=1}^kb
	_i(x+\delta)^{s_i}\right)-x$ and $\tau^{-1}(x)$ is the compositional inverse of $\tau(x)$ over $\gf_q.$ 
	Hence, 
	$$P_1^{-1}(x)+x=(P_2^{-1}(x)+x)^q.$$
	We are done. 
\end{proof}
\begin{remark} By Proposition \ref{th11}, we will only select either $P_1(x)$ or  $P_2(x)$ for our study in the following subsections. 
	\end{remark}

According to  Lemma \ref{th1}, it can be observed  that obtaining the  compositional inverse of $\tau(x)$ over $\gf_q$ enables us to determine the compositional inverse of $P(x)$ over $\gf_{q^2}.$ Thus, we split the subsequent subsections based on the form of  $\tau(x)$, $\tau^2(x)$, or $\tau^4(x).$  In each subsection, we provide a detailed proof for one representative result and only present the equations of $\tau(x)$, $\tau^2(x)$, or $\tau^4(x)$
 for the other theorems.

\subsection{ The form of $\tau(x)$, $\tau^2(x)$ or $\tau^4(x)$ is $Ax^{2}+C$ or $Ax+B$
}\label{sec3.1}
	In following two results, we will give two classes of involution of the form $\sum_{i=1}^{d}b_i(x^q+x+\delta)^{s_i}-x$ over $\gf_{q^2}.$ The results can be obtained directly from Lemma \ref{th1}. We omit the details here. 
\begin{theorem}
	Let $q$ be a power of $2.$ For $i=1, \cdots, k, $ assume that $b_i\in \gf_q^*, \delta \in \gf_{q^2}$ and $s_i$ are positive integers with $s_iq\equiv s_i \pmod{q^2-1}$. Then the polynomial
	$$P(x)=\sum_{i=0}^kb_i(x^q+x+\delta)^{s_i}+x$$  is an involution over $\gf_{q^2}$.
\end{theorem}
\begin{theorem}
	Let $q$ be a power of $2.$ For $i=1, \cdots, k, $ assume that $b_i, \delta \in \gf_{q}^*$ and $s_i$ are positive integers. Then the polynomial
	$$P(x)=\sum_{i=0}^kb_i(x^q+x+\delta)^{s_i}+x$$  is an involution over $\gf_{q^2}$.
\end{theorem}
\begin{remark}
	Based on the above theorem, we only consider $\delta \notin \gf_q$ when  calculating  the compositional inverse of the permutation polynomials of the form  $P(x)=\sum_{i=0}^kb_i(x^q+x+\delta)^{s_i}+x$ over $\gf_{q^2}$ with  $q$ being a power of $2$ and $b_i \in \gf_{q}$ in subsections \ref{sec3.1} and \ref{sec3.2}.
\end{remark}

\begin{theorem}\label{th28}
	Let $q=2^m\geq4$ with a  positive integer $m$. Assume that  $ \delta \in \gf_{q^2}$  and $b\in \gf_q^*$ with $b^4\Tr_{q^2/q}(\delta)=1.$   Then the compositional inverse of  
	$$P(x)=b(x^q+x+\delta)^{q(2q+3)/4}+x$$ over  $\gf_{q^2}$ is $$P^{-1}(x)=x+b\left(\Tr_{q^2/q}(\delta)^{-1}(x^q+x)^2+\delta^{q+1}\Tr_{q^2/q}(\delta)^{-1}+\delta\right)^{q(2q+3)/4}.$$
\end{theorem}

\begin{proof}
	\cite{wu2023some} had shown that if $b^4\Tr_{q^2/q}(\delta)=1,$ then the polynomial $P(x)=b(x^q+x+\delta)^{q(2q+3)/4}+x$ is a permutation polynomial over  $ \gf_{q^2}.$ Therefore, by Lemma \ref{th1},  $\tau(x)=b(x+\delta)^{q^2(2q+3)/4}+b(x+\delta)^{q(2q+3)/4}+x$ permutes  $\gf_{q}. $
	Moreover,  since $b^4\Tr_{q^2/q}(\delta)=1,$  we have
	\begin{align*}
		\tau(x)^4=&\,b^4(x+\delta)^{2q+3}+b^4(x+\delta)^{3q+2}+x^4\\
		=&\, b^4\Tr_{q^2/q}(\delta)(x^2+\delta^2)(x^2+\delta^{2q})+x^4\\
		=&\, \Tr_{q^2/q}(\delta)^2x^2+\delta^{2q+2},			
	\end{align*}
	or \begin{equation*}
		\tau(x)=x^{q/4}\circ (\Tr_{q^2/q}(\delta)^2x+\delta^{2q+2})\circ x^2.
	\end{equation*}
	This yields 
	\begin{align*}
		\tau^{-1}(x)=&\, x^{q/2}\circ (\Tr_{q^2/q}(\delta)^{-2}x+\delta^{2q+2}\Tr_{q^2/q}(\delta)^{-2})\circ x^4\\
		=&\, \Tr_{q^2/q}(\delta)^{-1}x^2+\delta^{q+1}\Tr_{q^2/q}(\delta)^{-1}.
	\end{align*}
	Hence,  by Lemma \ref{th1}, the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\Tr_{q^2/q}(\delta)^{-1}(x^q+x)^2+\delta^{q+1}\Tr_{q^2/q}(\delta)^{-1}+\delta\right)^{q(2q+3)/4}.$$
	We are done. 
\end{proof}

Given $\tau(x)^2= \Tr_{q^2/q}(\delta)x+\delta^{q+1},$  we have the following result. 
\begin{theorem}
	Let $q$ be a power of $2.$ Assume that $b \in \gf_{q}^*$ and $ \delta\in \gf_{q^2}$ with $b\Tr_{q^2/q}(\delta)=1$. 	Then the compositional inverse of $$P(x)=b(x^q+x+\delta)^{\frac{q^2+q}{2}+1}+x$$  over $\gf_{q^2}$ is  
	\begin{align*}
		P^{-1}(x)=&\, x+b\left(\Tr_{q^2/q}(\delta)^{-1}(x^q+x)^2+\Tr_{q^2/q}(\delta)^{-1}\delta^{q+1}+\delta\right)^{1+(q^2+q)/2}.
	\end{align*}
\end{theorem}
Similarly, we have following theorem. We only give $\tau^2(x)=(\Tr_{q^2/q}(b)^2+1)x^2+\Tr_{q^2/q}(b)^2\Tr_{q^2/q}(\delta)x+\delta^{q+1}\Tr_{q^2/q}(b)^2$ here.

\begin{theorem}\label{tha4}	
	Let $q$ be a power of $2.$  Assume that $b\in \gf_{q^2}^*$ and $\delta \in \gf_{q^2}.$ Let the polynomial
	$$P(x)=b(x^q+x+\delta)^{q(q+1)/2}+x$$ permute $\gf_{q^2}.$ \\
	If $b\in \gf_q$ or $\delta \in \gf_q$, then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=b\left(\frac{x^q+x}{\Tr_{q^2/q}(b)+1}+\frac{\delta^{(q+1)q/2}\Tr_{q^2/q}(b)}{\Tr_{q^2/q}(b)+1}+\delta\right)^{q(q+1)/2}+x.$$	
	If $\Tr_{q^2/q}(b)=1, $ then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=b\left(\Tr_{q^2/q}(b)^{-1}(x^q+x)^2+\Tr_{q^2/q}(b)^{-1}+\delta\right)^{q(q+1)/2}+x.$$	
\end{theorem}
Given $\tau(x)=b\Tr_{q^2/q}(\delta)x^2+b\delta^{q+1}\Tr_{q^2/q}(\delta), $ we have the following theorem.

\begin{theorem}
	Let  $q$ be a power of $2.$ Assume that $b \in \gf_{q}^*$ and $ \delta \in \gf_{q^2}$ with $b\Tr_{q^2/q}(\delta)^2=1.$ Then 
	the compositional inverse of
	$$P(x)=b(x^q+x+\delta)^{q+2}+x$$ over $\gf_{q^2}$ is \begin{align*}
		P^{-1}(x)=&\,b\left(\left(\Tr_{q^2/q}(\delta)(x^q+x)+\delta^{q+1}\right)^{q/2}+\delta \right)^{q+2}+x.
	\end{align*}
\end{theorem}

Given  $\tau(x)= b\Tr_{q^2/q}(\delta)^{2^i}x^{2}+b\delta^{q+1}\Tr_{q^2/q}(\delta)^{2^i}$ here, we have following result.
\begin{theorem}\label{th24}
	Let $q=2^m$ with a  positive integer $m.$ Assume that  $ \delta \in  \gf_{q^2}$, $b\in \gf_{q}^*$ and $i$ is a non-negative  integer with $b\Tr_{q^2/q}(\delta)^{2^i+1}=1. $ 
	Then the compositional inverse of $P(x)=b(x^q+x+\delta)^{2^i+q+1}+x$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\left(\Tr_{q^2/q}(\delta)(x^q+x)+\delta^{q+1}\right)^{q/2}+\delta\right)^{2^i+q+1}.$$
\end{theorem}

G. Li and X. Cao \cite{li2023several} presented that the polynomial $f(x)=(x^{p^m}-x+\delta)^{l(p^m+1)+p^j}+(x^{p^m}-x+\delta)^{l(p^m+1)+p^{m+j}}+x$ is a permutation polynomial over $\gf_{p^{2m}},$ where $p$ is prime, and $l, j $ and $m$ are positive integers in Proposition 8. We improve this result and give the compositional inverse of permutation polynomials of this form in the following theorem.  

\begin{theorem}\label{thq2}
	Let $q$ be an odd prime power. For $1\leq i \leq k, $  $b_i\in \gf_q,$  $\delta\in \gf_{q^2},$ $l_i,$  $s_i $ are positive integers with $2\nmid l_i$. Let $\varepsilon$ be a primitive element of $\gf_{q^2}$ and $t=(q+1)/2.$ Then
	the polynomial 
	$$P(x)=-x+\sum_{i=1}^kb_i\varepsilon^{tl_i}\left((x^q+x+\delta)^{s_i}+(x^q+x+\delta)^{qs_i}\right)$$ 
	permutes $\gf_{q^2}$ and  the compositional inverse of  $P(x)$ over $\gf_{q^2}$ is  
	$$P(x)^{-1}=-x+\sum_{i=1}^kb_i\varepsilon^{tl_i}\left((-x^q-x+\delta)^{s_i}+(-x^q-x+\delta)^{qs_i}\right).$$
\end{theorem}

Similarly, we have the following result. We only give $\tau(x)=\delta\sum_{i=0}^k\Tr_{q^2/q}(b_i)+\left(\sum_{i=0}^k\Tr_{q^2/q}(b_i)-1\right)x$ here. 
\begin{theorem}\label{thq3}
	Let $q$ be a prime power.  For $1\leq i\leq k$, assume that $\delta \in \gf_q,$  $b_i \in \gf_{q^2}$ and $s_i$ are positive integers. Then the polynomial
	$$P(x)=\sum_{i=1}^kb_i(x^q+x+\delta)^{s_i(q-1)+1}-x$$
	permutes $\gf_{q^2}$ if and only if $\sum_{i=0}^k\Tr_{q^2/q}(b_i)\neq1$
	Moreover, if $P(x)$ permutes $\gf_{q^2}$, the compositional inverse of $P(x)$ over $\gf_{q^2}$ is
	$$P^{-1}(x)=\sum_{i=0}^kb_i\left(\frac{x^q+x}{\sum_{i=0}^k\Tr_{q^2/q}(b_i)-1}+\frac{\delta\sum_{i=0}^k\Tr_{q^2/q}(b_i)}{\sum_{i=0}^k\Tr_{q^2/q}(b_i)-1}+\delta\right)^{s_i(q-1)+1}-x.$$
\end{theorem}

\subsection{ The form of $\tau(x)$ or $\tau^4$ is $Ax^{4}+Bx^2+Cx+D $ 
}\label{sec3.2}

Throughout this subsection, let $q=2^m$ with a positive integer $m,$ $b\in\gf_{q}^*,$ and $\delta \in \gf_{q^2}\setminus \gf_q.$
\begin{theorem}
	If $q\geq 4$ and the polynomial
	$$P(x)=b(x^q+x+\delta)^{1+(q^2+q)/4}+x$$ permutes $\gf_{q^2},$ then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(S_{m-2-i}^{2^{i+1}}+D^{1-2^{i+1}}S_i\right)((x^q+x)^4+B)^{2^i}\right)^{1+(q^2+q)/4},$$	 	where
	$B=b^4(\delta^q+\delta)^4\delta^{q+1}, C=b^4(\delta^q+\delta)^4, D=b^4(\delta^q+\delta)^5$ and $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$
	
\end{theorem}

\begin{proof}
	
	Since the polynomial $P(x)=b(x^q+x+\delta)^{1+(q^2+q)/4}+x$ permutes $\gf_{q^2}$, we obtain  that
	$\tau(x)=b(x+\delta)^{q+(q^3+q^2)/4}+b(x+\delta)^{1+(q^2+q)/4}
	+x$  permutes $\gf_{q}$ by Lemma \ref{th1},  and the latter condition is equivalent to $\tau(x)^4$ is a permutation polynomial over  $\gf_{q}.$ 		Furthermore, we have 
	\begin{align}\label{1eq1+(q^2+q)/4}
		\tau(x)^4=&\,b^4(x+\delta)^{5q+1}+b^4(x+\delta)^{q+5}
		+x^4\nonumber\\
		=&\, b^4(x+\delta)^{q+1}\left((x+\delta)^{4q}+(x+\delta)^{4}\right)+x^4\nonumber\\
		=&\,x^4+b^4(\delta^q+\delta)^4x^2+b^4(\delta^q+\delta)^5x
		+b^4(\delta^q+\delta)^4\delta^{q+1}
	\end{align}	
	Put $B=b^4(\delta^q+\delta)^4\delta^{q+1}, C=b^4(\delta^q+\delta)^4, D=b^4(\delta^q+\delta)^5.$
	Then by \eqref{1eq1+(q^2+q)/4}, we have
	\begin{align*}
		\tau(x)^4
		=&\, (x+ B)\circ	(x^4+Cx^2+Dx),
	\end{align*}
	or \begin{equation} \label{1eq1+(q
			^2+q)/4}
		\tau(x)=x^{q/4}\circ(x+ B)\circ	(x^4+Cx^2+Dx).
	\end{equation}
	Since  $\tau(x)$ permutes  $\gf_{q},$ it follows from Lemma \ref{421} and \eqref{1eq1+(q
		^2+q)/4} that the compositional inverse of $\tau(x)$ over $\gf_q$ is 
	\begin{align}\label{2eq1+(q
			^2+q)/4}
		\tau^{-1}(x)=&\,(x^4+Cx^2+Dx)^{-1}\circ (x+ B)^{-1}\circ x^4\nonumber\\
		=&\,\sum_{i=0}^{m-1}\left(S_{m-2-i}^{2^{i+1}}+D^{1-2^{i+1}}S_i\right)x^{2^i}\circ (x+B)\circ x^4\nonumber\\
		=&\,\sum_{i=0}^{m-1}\left(S_{m-2-i}^{2^{i+1}}+D^{1-2^{i+1}}S_i\right)(x^4+B)^{2^i},
	\end{align}
	where $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$
	Lemma \ref{th1} and \eqref{2eq1+(q
		^2+q)/4}  implies that the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(S_{m-2-i}^{2^{i+1}}+D^{1-2^{i+1}}S_i\right)((x^q+x)^4+B)^{2^i}\right)^{1+(q^2+q)/4}.$$	
	We are done. 		
\end{proof}
Similarly, we have the following theorem.  We only give $\tau^4(x)=(x+b^4\delta^{q+1}\Tr_{q^2/q}(\delta))\circ (x^4+b^4\Tr_{q^2/q}(\delta)x^2+b^4\Tr_{q^2/q}(\delta)^2x)$ here. 

\begin{theorem}\label{th27}
	If $q\geq 4$ and the polynomial
	$$P(x)=b(x^q+x+\delta)^{q(2q+1)/4}+x$$ permutes $\gf_{q^2}$, then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(S_{m-2-i}^{2^{i+1}}+D^{1-2^{i+1}}S_i\right)\left(x^{4q}+x^4+B\right)^{2^i}\right)^{q(2q+1)/4},$$
	where	 $B=b^4\delta^{q+1}\Tr_{q^2/q}(\delta),$  $C=b^4\Tr_{q^2/q}(\delta),$ $D= b^4\Tr_{q^2/q}(\delta)^2,$ and $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$
\end{theorem}

Similarly, we have the following result. We only give $\tau(x)=b\Tr_{q^2/q}(\delta)x^4+ b\Tr_{q^2/q}(\delta)^3x^2
+x+  b\delta^{2q+2}\Tr_{q^2/q}(\delta)$ here. 
\begin{theorem}
	If the polynomial
	$P(x)=b(x^q+x+\delta)^{2q+3}+x$ permutes $\gf_{q^2}$, then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(D^{2^i}S_{m-2-i}^{2^{i+1}}+D^{1-2^{i}}S_i\right)\left( x^q+x+B\right)^{2^i}\right)^{2q+3},$$
	where $A= b\Tr_{q^2/q}(\delta), B=b\delta^{2q+2}\Tr_{q^2/q}(\delta), C=\Tr_{q^2/q}(\delta)^2, D=1/A,$ and $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$

\end{theorem}

Similarly, we have the following result. We only give $\tau(x)=(b\Tr_{q^2/q}(\delta)^{2^i}x+b\Tr_{q^2/q}(\delta)^{2^i}\delta^{2q+2})\circ (x^4+\Tr_{q^2/q}(\delta)^2x^2+b^{-1}\Tr_{q^2/q}(\delta)^{-2^i}x)$ here.

\begin{theorem}\label{th25}
	For a non-negative integer $i$, if the polynomial
	$$P(x)=b(x^q+x+\delta)^{2q+2^i+2}+x$$ permutes $\gf_{q^2},$ then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(D^{2^i}S_{m-2-i}^{2^{i+1}}+D^{1-2^{i}}S_i\right)\left( x^q+x+B\right)^{2^i}\right)^{2q+2^i+2},$$
	where $A= b\Tr_{q^2/q}(\delta)^{2^i}, B=b\Tr_{q^2/q}(\delta)^{2^i}\delta^{2q+2}, C=\Tr_{q^2/q}(\delta)^2, D=1/A,$ and $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$	
\end{theorem}

Similarly, we have the following result. We only give $
\tau(x)=(b\Tr_{q^2/q}(\delta)^2x+b\delta^{6q}+b\delta^6)\circ	(x^4+\Tr_{q^2/q}(\delta)^2x^2+b^{-1}\Tr_{q^2/q}(\delta)^{-2}x).		
$

\begin{theorem}\label{th29}
	If the polynomial
	$$P(x)=b(x^q+x+\delta)^{6}+x$$ permutes $\gf_{q^2},$ then the compositional inverses of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=x+b\left(\delta+\sum_{i=0}^{m-1}\left(D^{2^i}S_{m-2-i}^{2^{i+1}}+D^{1-2^{i}}S_i\right)\left( x^q+x+B\right)^{2^i}\right)^{6},$$
	where $A=b\Tr_{q^2/q}(\delta)^2,$ $B=b\delta^{6q}+b\delta^6,$ $C=\Tr_{q^2/q}(\delta)^2,$ $D=A^{-1}$ and  $S_i$ is a sequence with $S_{-1}=0,$ $S_0=1, $ $S_i=C^{2^{i-1}}S_{i-1}+D^{2^{i-1}}S_{i-2}.$
\end{theorem}

\subsection{ The form of $\tau(x)$ is $Ax^{p^i}+Bx+C (p^i>2)$
 }

\begin{theorem}\label{thq1}
	Let $q=p^m$ with an odd prime $p$ and a positive integer $m.$  Assume that $\delta \in \gf_{q^2}$, $b_1, b_2, b_3 \in \gf_q,$ and  $\varepsilon \in \gf_{q^2}$ is a primitive element of $\gf_{q^2}.$ For a non-negative integer $i$ with $i< 2m$, let $j,d $ be a non-negative integers with $j<m$,  $j\equiv i \pmod{m},$ and $d=\gcd(m, j).$ For odd integers $d_1, d_2, d_3$ and $t=(q+1)/2,$
let $$P(x)=b_1\varepsilon^{td_1}(x^q+x+\delta)^{p^i+q}+b_2\varepsilon^{td_2}(x^q+x+\delta)^{p^i+1}+b_3\varepsilon^{td_3}(x^q+x+\delta)^{2p^i}-x$$ be a  polynomial over $\gf_{q^2}.$ \\	
	If $i=0, m$, then $P(x)$ is a permutation polynomial over $\gf_{q^2}$ if and only if $A-B\neq0.$ Moreover, if the latter condition holds,  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
\begin{align*}P^{-1}(x)=&\,b_1\varepsilon^{td_1}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{p^i+q}\\
	&\,+b_2\varepsilon^{td_2}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{p^i+1}\\
	&\,+b_3\varepsilon^{td_3}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{2p^i}-x.
\end{align*}
If $i \neq 0, m,$  $A=0, $ and $B\neq0,$  then $P(x)$ is a permutation  polynomial over $\gf_{q^2}$ and  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
\begin{align*}P^{-1}(x)=&\,b_1\varepsilon^{td_1}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{p^i+q}\\
	&\,+b_2\varepsilon^{td_2}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{p^i+1}\\
	&\,+b_3\varepsilon^{td_3}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{2p^i}-x.
\end{align*}
If $i \neq 0, m,$  $A\neq0, $ and $B=0,$ then $P(x)$ is a permutation  polynomial over $\gf_{q^2}$ and   the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
\begin{align*}
	P^{-1}(x)=&\, b_1\varepsilon^{td_1}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{p^i+q}\\
	&\,+b_2\varepsilon^{td_2}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{p^i+1}\\
	&\,+b_3\varepsilon^{td_3}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{2p^i}-x.
\end{align*}
If  $i \neq 0, m$ and  $AB\neq0, $ then $P(x)$ permutes $\gf_{q^2}$ if and only if $N_{p^m/p^d}(B/A)\neq1$. Moreover, if  $P(x)$ permutes $\gf_{q^2}$,  then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 	\begin{align*}
	P^{-1}(x)=&\, b_1\varepsilon^{td_1}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{p^i+q}\\	
	+&\,b_2\varepsilon^{td_2}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{p^i+1}\\
	+&\,b_3\varepsilon^{td_3}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{2p^i}\\-&\,x.
\end{align*}
Here, \begin{align*} A=&\,(\delta^q-\delta)(b_1\varepsilon^{td_1}-b_2\varepsilon^{td_2})+2b_3\varepsilon^{td_3}(\delta^{p^i}-\delta^{p^iq}),\\ B=&\,1-(\delta^{p^iq}-\delta^{p^i})(b_1\varepsilon^{td_1}+b_2\varepsilon^{td_2}),\\  C=&\,(\delta^{p^i+q}-\delta^{p^iq+1})b_1\varepsilon^{td_1}+(\delta^{p^i+1}-\delta^{p^iq+q})b_2\varepsilon^{td_2}+(\delta^{2p^i}-\delta^{2p^iq})b_3\varepsilon^{td_3}.
\end{align*}
\end{theorem}

\begin{proof}
	Since $\varepsilon \in \gf_{q^2}$ is a primitive element of $\gf_{q^2}$ and $t=(q+1)/2$, we have $\varepsilon^{(q-1)t}=-1,$ or 
	\begin{equation}\label{a1}
		\varepsilon^{qt}=-\varepsilon^t
	\end{equation}
	Lemma \ref{th1} yields that $P(x)$ permutes $\gf_{q^2}$
	if and only if $\tau(x)=b_1\varepsilon^{qtd_1}(x+\delta)^{p^iq+1}+b_1\varepsilon^{td_1}(x+\delta)^{p^i+q}+b_2\varepsilon^{qtd_2}(x+\delta)^{p^iq+q}+b_2\varepsilon^{td_2}(x+\delta)^{p^i+1}+b_3\varepsilon^{qtd_3}(x+\delta)^{2p^iq}+b_3\varepsilon^{td_3}(x+\delta)^{2p^i}-x.$ Note that $\tau(x)$ is a polynomial over $\gf_q.$
	Moreover, by \eqref{a1},  we have
	{\small	\begin{align}\label{1eqq+12q}
			\tau(x)=&\,b_1\varepsilon^{qtd_1}(x+\delta)^{p^iq+1}+b_1\varepsilon^{td_1}(x+\delta)^{p^i+q}+b_2\varepsilon^{qtd_2}(x+\delta)^{p^iq+q}\nonumber\\
			&\,+	b_2\varepsilon^{td_2}(x+\delta)^{p^i+1}+b_3\varepsilon^{qtd_3}(x+\delta)^{2p^iq}+b_3\varepsilon^{td_3}(x+\delta)^{2p^i}-x\nonumber\\
			=&\, -b_1\varepsilon^{td_1}(x^{p^i+1}+\delta x^{p^i}+\delta^{p^iq}x+\delta^{p^iq+1})+b_1\varepsilon^{td_1}(x^{p^i+1}+\delta^q x^{p^i}+\delta^{p^i}x+\delta^{p^i+q})\nonumber\\
		&\,-b_2\varepsilon^{td_2}(x^{p^i+1}+\delta^q x^{p^i}+\delta^{p^iq}x+\delta^{p^iq+q})+b_2\varepsilon^{td_2}(x^{p^i+1}+\delta x^{p^i}+\delta^{p^i}x+\delta^{p^i+1})\nonumber\\
		&\,	-b_3\varepsilon^{td_3}(x^{2p^i}+2\delta^{p^iq} x^{p^i}+\delta^{2p^iq})+b_3\varepsilon^{td_3}(x^{2p^i}+2\delta^{p^i} x^{p^i}+\delta^{2p^i})-x\nonumber\\
			=&\, Ax^{p^j}-Bx+C \end{align}}
	where $j$ is a non-negative integer with $j<m$ and $j\equiv i \pmod{m}$ and
	\begin{align*} A=&\,(\delta^q-\delta)(b_1\varepsilon^{td_1}-b_2\varepsilon^{td_2})+2b_3\varepsilon^{td_3}(\delta^{p^i}-\delta^{p^iq}),\\ B=&\,1-(\delta^{p^iq}-\delta^{p^i})(b_1\varepsilon^{td_1}+b_2\varepsilon^{td_2}),\\  C=&\,(\delta^{p^i+q}-\delta^{p^iq+1})b_1\varepsilon^{td_1}+(\delta^{p^i+1}-\delta^{p^iq+q})b_2\varepsilon^{td_2}+(\delta^{2p^i}-\delta^{2p^iq}-)b_3\varepsilon^{td_3}.
	\end{align*}
	
	If $i=0, m$, then $\tau(x)=(A-B)x+C$ by \eqref{1eqq+12q}, so $\tau(x)$ permutes $\gf_q$ if and only if $A-B\neq0.$ Moreover, if the latter condition holds, then  $\tau^{-1}(x)=(A-B)^{-1}x-(A-B)^{-1}C$ by \eqref{1eqq+12q} and $P(x)$ permutes $\gf_{q^2}$. Hence, the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}P^{-1}(x)=&\,b_1\varepsilon^{td_1}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{p^i+q}\\
		&\,+b_2\varepsilon^{td_2}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{p^i+1}\\
		&\,+b_3\varepsilon^{td_3}((A-B)^{-1}(x^q+x)-(A-B)^{-1}C+\delta)^{2p^i}-x.
		\end{align*}

	Next, we will do the case  $i\neq 0, m.$
	Notice that $\tau(x)$ is a affine polynomial over $\gf_q.$ Then $\tau(x)$ permutes $\gf_q$ if and only if $A=0$, or $B=0, $ or  $N_{p^m/p^d}(B/A)\neq1$ by Lemma \ref{binomial}, where $d=\gcd(m, j).$ 
	
	Moreover, if $A=0$ and $B\neq0$, then by \eqref{1eqq+12q}, we get  $\tau(x)=-Bx+C,$ and so $\tau^{-1}(x)=-B^{-1}x+B^{-1}C.$ This yields that  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
			\begin{align*}P^{-1}(x)=&\,b_1\varepsilon^{td_1}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{p^i+q}\\
		&\,+b_2\varepsilon^{td_2}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{p^i+1}\\
		&\,+b_3\varepsilon^{td_3}(-B^{-1}(x^q+x)+B^{-1}C+\delta)^{2p^i}-x.
	\end{align*}
	by Lemma \ref{th1}.
	
	If $A\neq0$ and $B= 0$, then by \eqref{1eqq+12q}, we obtain $\tau(x)=Ax^{p^j}+C,$ and so $\tau^{-1}(x)=
	A^{-p^{m-j}}x^{p^{m-j}}-(C/A)^{p^{m-j}}.$
	Consequently, by Lemma \ref{th1}, the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}
		P^{-1}(x)=&\, b_1\varepsilon^{td_1}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{p^i+q}\\
		&\,+b_2\varepsilon^{td_2}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{p^i+1}\\
		&\,+b_3\varepsilon^{td_3}(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}-(C/A)^{p^{m-j}}+\delta)^{2p^i}-x.
	\end{align*}
	
	If  $AB\neq0$ and $N_{p^m/p^d}(B/A)\neq1$,  then $\tau(x)=Ax^{p^j}-Bx+C=(Ax+C)\circ (x^{p^j}-Bx/A)$ permutes $\gf_q$, and so,  by Lemma \ref{binomial}, we have 
	\begin{align*}
		\tau^{-1}(x)=&\,\left(\frac{N_{p^m/p^d}(B/A)}{1-N_{p^m/p^d}(B/A)}\sum_{k=0}^{m/d-1}(B/A)^{-\frac{p^{(k+1)j}-1}{p^j-1}}x^{p^{kj}}\right)\circ (A^{-1}x-A^{-1}C)\\
		=&\,\frac{N_{p^m/p^d}(B/A)}{1-N_{p^m/p^d}(B/A)}\sum_{k=0}^{m/d-1}(A/B)^{\frac{p^{(k+1)j}-1}{p^j-1}}(A^{-1}x-A^{-1}C)^{p^{kj}}.
	\end{align*}
	
	Hence, it follows from Lemma \ref{th1} that  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}
	P^{-1}(x)=&\, b_1\varepsilon^{td_1}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{p^i+q}\\	
	+&\,b_2\varepsilon^{td_2}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{p^i+1}\\
+&\,b_3\varepsilon^{td_3}\left(\delta+\frac{N_{p^m/p^d}(\frac{B}{A})}{1-N_{p^m/p^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}-\frac{C}{A}\right)^{p^{kj}}\right)^{2p^i}\\-&\,x.
	\end{align*}
	This completes the proof.
\end{proof}
Similarly, we have following result. We omit the details.
\begin{theorem}\label{thq1}
	Let $q=2^m$ with  a positive integer $m.$  Assume that $\delta \in \gf_{q^2}$, $b_1, b_2 \in \gf_q,$ and $b_3 \in \gf_{q^2}. $  For a non-negative integer $i$ with $i< 2m$, let  $j, d$ be a non-negative integers with $j<m, $  $j\equiv i \pmod{m},$ and $d=\gcd(m, j).$ 	
And	let $$P(x)=b_1(x^q+x+\delta)^{2^i+q}+b_2(x^q+x+\delta)^{2^i+1}+b_3(x^q+x+\delta)^{2^i}+x$$ be a  polynomial over $\gf_{q^2}.$ \\
	If $i=0, m$, then $P(x)$ is a permutation polynomial over $\gf_{q^2}$ if and only if $A+B\neq0.$ Moreover, if the latter condition holds,  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}P^{-1}(x)=&\,b_1((A+B)^{-1}(x^q+x)+(A+B)^{-1}C+\delta)^{2^i+q}\\
		&\,+b_2((A+B)^{-1}(x^q+x)+(A+B)^{-1}C+\delta)^{p^i+1}\\
		&\,+b_3((A+B)^{-1}(x^q+x)+(A+B)^{-1}C+\delta)^{2^i}+x.
	\end{align*}
	If $i \neq 0, m,$  $A=0, $ and $B\neq0,$  then $P(x)$ is a permutation  polynomial over $\gf_{q^2}$ and  the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}P^{-1}(x)=&\,b_1(B^{-1}(x^q+x)+B^{-1}C+\delta)^{2^i+q}\\
		&\,+b_2(B^{-1}(x^q+x)+B^{-1}C+\delta)^{2^i+1}\\
		&\,+b_3(B^{-1}(x^q+x)+B^{-1}C+\delta)^{2^i}+x.
	\end{align*}	
	If $i \neq 0, m,$  $A\neq0, $ and $B=0,$ then $P(x)$ is a permutation  polynomial over $\gf_{q^2}$ and the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	\begin{align*}
		P^{-1}(x)=&\, b_1(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}+(C/A)^{p^{m-j}}+\delta)^{2^i+q}\\
		&\,+b_2(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}+(C/A)^{p^{m-j}}+\delta)^{p^i+1}\\
		&\,+b_3(A^{-p^{m-j}}(x^q+x)^{p^{m-j}}+(C/A)^{p^{m-j}}+\delta)^{2^i}+x.
	\end{align*}	
	If  $i \neq 0, m$ and  $AB\neq0, $ then $P(x)$ permutes $\gf_{q^2}$ if and only if $N_{2^m/2^d}(B/A)\neq1$. Moreover, if  $P(x)$ permutes $\gf_{q^2}$,  then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 	\begin{align*}
		P^{-1}(x)=&\, b_1\left(\delta+\frac{N_{2^m/2^d}(\frac{B}{A})}{1+N_{2^m/2^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}+\frac{C}{A}\right)^{2^{kj}}\right)^{2^i+q}\\	
		+&\,b_2\left(\delta+\frac{N_{2^m/2^d}(\frac{B}{A})}{1+N_{2^m/2^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}+\frac{C}{A}\right)^{p^{kj}}\right)^{2^i+1}\\
		+&\,b_3\left(\delta+\frac{N_{2^m/2^d}(\frac{B}{A})}{1+N_{2^m/2^d}\left(\frac{B}{A}\right)}\sum_{k=0}^{m/d-1}\left(\frac{A}{B}\right)^{\frac{p^{(k+1)j}-1}{p^j-1}}\left(\frac{x^q+x}{A}+\frac{C}{A}\right)^{p^{kj}}\right)^{2^i}+x.
	\end{align*}
	Here, \begin{align*} A=&\,(\delta+\delta^q)(b_1+b_2)+b_3+b_3^q,\\ B=&\,1+(\delta^{2^iq}+\delta^{2^i})(b_1+b_2),\\  C=&\,b_1(\delta^{2^iq+1}+\delta^{2^i+q})+b_2(\delta^{2^iq+q}+\delta^{2^i+1})+b_3^q\delta^{2^iq}+b_3a\delta^{2^i}.
	\end{align*}
\end{theorem}

Similarly, we also have following result. We only give $
	\tau(x)= -bx^3+\left(b(\delta^q-\delta)^2-1\right)x+b\delta^{q+1}(\delta^q+\delta)$ here.

\begin{theorem}\label{3 q+2th}
	Let $q=3^m$ with a  positive integer $m.$   Assume that $b \in \gf_{q}^*$ and $ \delta \in \gf_{q^2}.$  Let $A=-b, B=b\delta^{q+1}(\delta^q+\delta), C=(\delta^q-\delta)^2-b^{-1}.$
	And let
	$$P(x)=b(x^q+x+\delta)^{q+2}-x$$ permute $\gf_{q^2}.$\\	
	If $(\delta^q-\delta)^2-b^{-1}=0,$ then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is  	$$P^{-1}(x)=b\left(A^{-q/3}((x^q+x)-B)^{q/3}+\delta\right)^{q+2}-x.$$\\	
	If	$(\delta^q-\delta)^2-b^{-1}$ is not a square in $\gf_q,$ then the compositional inverse of $P(x)$ over $\gf_{q^2}$ is 
	$$P^{-1}(x)=b\left(\delta+\frac{N_{3^m/3}(C)}{1-N_{3^m/3}(C)}\sum_{i=0}^{m-1}A^{-3^i}C^{-\frac{3^{i+1}-1}{2}}(x^q+x-B)^{3^i}\right)^{q+2}-x.$$
	
\end{theorem}





\noindent
{\bf Acknowledgments}
\\

 P. Yuan was supported by the National Natural Science Foundation of China (Grant No. 12171163), Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515010589). D. Wu was supported by Guangdong Basic and Applied Basic Research Foundation (Grant No. 2020A1515111090).
\\

\end{document}